Suppose a perfectly competitive firm faces the following situation: P = $8, output = 2,000, ATC = $6.50, and MC = $7. Which statement is an accurate description of the firm's situation?

a.The firm incurs profits but is not maximizing its profits.

b.The firm is maximizing profits.

c.The firm incurs losses and is minimizing its losses.

d.The firm incurs losses but is not minimizing its losses.
Answer #1

Answer

Option a

a.The firm incurs profits but is not maximizing its profits.

Profit=(P-ATC)*Q

=(8-6.5)*2000=3000

but the profit is maximum at MR=MC so the firm needs to increase output up to MR=MC.